OpenStudy (anonymous):

Does anyone know how to do this? (I am either doing it wrong or doing my math incorrectly)

OpenStudy (anonymous):

i know how to do this

OpenStudy (anonymous):

ok so, numbers are 4,13,5,6,9 If the mean of these data is approximately 7, what is the population standard deviation for these data? (Round the answer to the nearest tenth.)

OpenStudy (anonymous):

The population standard deviation is defined as: $$s = \sqrt{\frac 1 {N-1} \sum_{i=1}^N (x_i - \bar x)^2}$$

OpenStudy (anonymous):

Ah, no, nevermind. 'Tis N, not N-1.

OpenStudy (anonymous):

oh uh, sorry, but do you know what we plug into that equation?

OpenStudy (anonymous):

N is the number of data values. Okay, so the simplified thing is: #1.) Take the mean. We have that; it's 7. We have that N = 5. #2.) Now we subtract the mean from each datum and square the result. Let's see: 4 - 7 = -3, 3^2 = 9 13 - 7 = 6, 6^2 = 36 5 - 7 = -2, 2^2 = 4 6 - 7 = -1, 1^1 = 1 9 - 7 = 2, 2^2 = 4 #3.) Now we take the mean of those values. 54/5, or 11.02. #4.) Finally, we take the square root of the final product. sqrt(11.02) = ?

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

fail, that's what I forgot to do, take the square root, wow

OpenStudy (anonymous):

thank-you so much